HETAV PATEL

Grade 11,

a proton enters a magneticfield of flux densitsy 1.5T with a velocity of 20x10 7 ms at an angle 30 with the field the force on the proton is

a proton enters a magneticfield of flux densitsy 1.5T with a velocity of 20x107 ms at an angle 30 with the field the force on the proton is

Grade:12

1 Answers

Kriti Agrawal
7 years ago
Force in a magnetic field equals the cross product of velocity and B vector multiplied by charge. Flux density is the magnetic field. Apply the formula F=q (v×B) which is qvb sin30. q=1.6×10^-19. You will straightaway arrive at the answer.